OpenStudy (anonymous):

integral sin(x)/3x find the first five non zero term centered around x=0. what is radius of convergence?

OpenStudy (anonymous):

it is sin(x)/3x and I need to find radius of convergence. I would really appreciate if anyone helps me out with this one.

OpenStudy (experimentx):

lin n->inf 1/(1/(2i+1)!)^(1/n)

OpenStudy (anonymous):

do you mean lim ?

OpenStudy (experimentx):

limit

OpenStudy (anonymous):

wouldnt answer be just zero?

OpenStudy (experimentx):

i found excellent example

OpenStudy (anonymous):

or do I have to use ratio or root test? The function we are taking limit is derived from taylor series??

OpenStudy (anonymous):

I was trying to list taylor series and try to convert the series into sigma of a function and take ratio or root test to find radius.

OpenStudy (experimentx):

ratio test i guess

OpenStudy (anonymous):

lin n->inf 1/(1/(2i+1)!)^(1/n) just to ratio test for this one

OpenStudy (experimentx):

http://tutorial.math.lamar.edu/Classes/CalcII/PowerSeries.aspx example 1 below ... BRB

OpenStudy (anonymous):

Thank you so much. I think my concern was how you got inf 1/(1/(2i+1)!)^(1/n) from sinx/3x

OpenStudy (anonymous):

how about 3x?

OpenStudy (experimentx):

divide x's by x ... 1/3 is just a constant.

OpenStudy (anonymous):

and therefore we get from sinx/3x to 1/(1/(2i+1)!)^(1/n)

OpenStudy (experimentx):

no ... that would be the nth root of of coefficint. plus 1/3 is missing.

OpenStudy (anonymous):

sorry but do you mean? then (1/(1/(2i+1)!)^(1/n)) +(1/3)

OpenStudy (experimentx):

radius of convergence is also expained here. http://blogs.ubc.ca/infiniteseriesmodule/units/unit-3-power-series/taylor-series/maclaurin-expansion-of-sinx/ you just have to write 2k for 2k+1 an have 1/3 out

OpenStudy (experimentx):

I'll be back in an hour.

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

I think I got stuck in the process. Kind of lost.

OpenStudy (anonymous):

can anyone help me out?

OpenStudy (experimentx):

i guess it would be same ... 0

OpenStudy (anonymous):

what do you mean??

OpenStudy (anonymous):

Here are the first few terms \[-\frac{x^{11}}{1317254400}+\frac {x^9}{9797760}-\frac{x^7}{105 840}+\frac{x^5}{1800}-\frac{x ^3}{54}+\frac{x}{3} \] Read it backwards

OpenStudy (anonymous):

\[ \sin(x) = \sum_{n=0}^{\infty} (-1)^{n+1}\frac { x^{2n+1}}{(2n+1)!} \] \[ \frac {\sin(x)}{3x} = \sum_{n=0}^{\infty} (-1)^{n+1}\frac { x^{2n}}{3(2n+1)!} \] \[ \int \frac {\sin(x)dx}{3x} = \sum_{n=0}^{\infty} (-1)^{n+1}\int \frac { x^{2n+1}dx}{3(2n+1)!}= \sum_{n=0}^{\infty} (-1)^{n+1 } \frac { x^{2n}dx}{3(2n+1)(2n+1)!} \] Apply the ratio test with \[ a_n = \frac { x^{2n}}{3(2n+1)(2n+1)!} \] to deduce that \[ R = \infty\]

OpenStudy (anonymous):

thank you !!